# Rapid formation of a very massive star $>\!\!50000\,\mathrm{M}_{\odot}$ and subsequently an IMBH from runaway collisions

### v1

@@ 4.2 N-body results: VMS evolution and IMBH formation
The hydrogen burning phase is set to end in 2.92.92.9\,2.9Myr, according to the SSE/BSE prescription Hurley et al. (2000a, 2002a, 2005). After the VMS reaches this point, it should start burning helium. At 4.674.674.67\,4.67Myr the VMS has a mass of $5.7362\times 10^{4}\leavevmode\nobreak\ \rm M_{\odot}$, with a radius of $1.9\times 10^{5}\rm R_{\odot}$ and effective age of 2.912.912.91\,2.91Myr. At this stage, the VMS becomes a helium main sequence (HeMS; or type 7 in SSE/BSE by Hurley et al. (2000a, 2002a, 2005))) star with a radius of $171\rm R_{\odot}$. During this helium phase, the VMS experiences three collisions with MS stars within 0.020.020.02\,0.02Myr before naturally collapsing into an IMBH. This process may result in mass loss of about $10\leavevmode\nobreak\ \%$ due to neutrino emission (Fryer et al., 2012; Kamlah et al., 2022a), leading to a mass of 51 $51\,625.99\leavevmode\nobreak\ \rm M_{\odot}$ at time 4.694.694.69\,4.69Myr. Finally, the IMBH undergoes additional collisions, accreting $50\%$ of the stellar companion (Rizzuto et al., 2021), reaching a final mass of 51 $51\,661.21\leavevmode\nobreak\ \rm M_{\odot}$ by the end of the simulation.
@@ 4.3 MOCCA results: VMS evolution and IMBH formation
In the MOCCA simulation, the star that eventually grows into a VMS has an initial mass of $\rm 130.6\ \leavevmode\nobreak\ \rm M_{\odot}$ and is initially located at a distance of 0.088 pc from the center of the cluster. Within 0.15 Myr, this massive star undergoes four collisions with low mass stars, slightly increasing its mass to approximately $\rm 132\ \leavevmode\nobreak\ \rm M_{\odot}$. For a brief period between about $0.18$ Myr and $0.21$ Myr, the VMS is part of a binary system. At $0.182$ Myr, the VMS first forms a binary with a $\rm 72.4\,M_{\odot}$ star due to 3BBF. Shortly after, at $0.183$ Myr, this companion is replaced by a $\rm 124\,M_{\odot}$ star in an exchange encounter. Following several flyby interactions, the VMS and its binary companion merge during a binary-single interaction at $0.186$ Myr. The VMS subsequently forms another binary, first with a $\rm 72.4\,M_{\odot}$ star, then with a $\rm 68\,M_{\odot}$ star, and later with a $\rm 174\,M_{\odot}$ star. The latter also merges with the VMS in a binary-single interaction at around $0.21$ Myr, following several flyby encounters. At this stage, the VMS is located at a distance of approximately $3\times 10^{-4}\ \rm pc$ from the cluster center. By $0.212$ Myr, its mass exceeds $\rm 500\,M_{\odot}$ following a collision with a $\rm 121.8\,M_{\odot}$ star, with which it was also briefly in a binary system. As described in Section 3.2, from this point onward, we disable 3BBF for the VMS. Its mass continues to increase rapidly as it undergoes runaway collisions with other MS stars in the cluster center. By 0.23 Myr, the VMS mass reaches $\rm 1000\ \leavevmode\nobreak\ \rm M_{\odot}$, and by 0.34 Myr, it exceeds $\rm 5000\ \leavevmode\nobreak\ \rm M_{\odot}$. The growth of the VMS in MOCCA is shown in the top-right panel of Fig. 8.
@@
During these early stages of VMS growth, collisions with other VMSs also strongly rejuvenate it. At 0.34 Myr, when its mass reaches $\rm 5000\ \leavevmode\nobreak\ \rm M_{\odot}$, its effective age is only 0.068 Myr, and its radius is approximately $\rm 68\ \mathrm{R}_{\odot}$. By 0.51 Myr, the VMS mass surpasses $\rm 10^{4}\ \leavevmode\nobreak\ \rm M_{\odot}$, at which point its effective age is 0.17 Myr, and its radius is $\rm 71.5\ \mathrm{R}_{\odot}$. The VMS continues its rapid growth, reaching $\rm 2.1\times 10^{4}\ \leavevmode\nobreak\ \rm M_{\odot}$ by 1 Myr. By this time, the cumulative number of collisions involving the VMS is approximately 3,815 (see Fig. 3). As the mass ratio of these collisions decreases with increasing VMS mass, its effective rejuvenation per collision also weakens. At 1 Myr, its effective age is 0.468 Myr. The evolution of the VMS’s effective age on the MS is shown in the bottom-right panel of Fig. 8.
@@
By 2 Myr, the VMS reaches a mass of $\rm 3.56\times 10^{4}\ \leavevmode\nobreak\ \rm M_{\odot}$, with an effective age of 1.1 Myr and a radius of $\rm 131\ \mathrm{R}_{\odot}$. Up until this point, the VMS has undergone more than 6,793 collisions. Between 2 Myr and 3 Myr, its growth slows, and by 3 Myr, it reaches a mass of $\rm 4.59\times 10^{4}\ \leavevmode\nobreak\ \rm M_{\odot}$ with an effective age of 1.77 Myr. This corresponds to approximately 61% of the MS lifetime of the VMS, which is set to 2.91 Myr in the SSE/BSE prescriptions used in NBODY6++GPU and MOCCA. As the VMS ages, its radius increases, reaching $\rm 523\ \mathrm{R}_{\odot}$ at 3 Myr. The evolution of the VMS’s radius is shown in the middle-right panel of Fig. 8.
@@
At 4.477 Myr, the VMS reaches a mass of $\rm 5.38\times 10^{4}\ \leavevmode\nobreak\ \rm M_{\odot}$, with a radius of $\rm 1.9\times 10^{5}\mathrm{R}_{\odot}$ and an effective age of 2.91 Myr, marking the MS turn-off time for the VMS. By this point, it has undergone 10,322 collisions. After this stage, the VMS evolves into a helium main sequence star (HeMS; type 7 in SSE/BSE) with a radius of $\rm 165\ \mathrm{R}_{\odot}$. Between 4.477 Myr and 4.485 Myr, the HeMS VMS undergoes approximately 15 additional collisions with lower-mass stars, reaching a final mass of $\rm 5.384\times 10^{4}\ \leavevmode\nobreak\ \rm M_{\odot}$. As the mass ratios in these collisions are extremely small, no rejuvenation is applied. At 4.485 Myr, the VMS evolves into an IMBH with a mass of $\rm 4.845\times 10^{4}\ \leavevmode\nobreak\ \rm M_{\odot}$. This IMBH then continues to grow through collisions, mostly with other stars, assuming that 50% of the stellar mass is accreted by the IMBH in each collision. By 5 Myr, it reaches a mass of $\rm 4.984\times 10^{4}\ \leavevmode\nobreak\ \rm M_{\odot}$.
@@
The short duration of the HeMS phase in the evolution of the VMS is a result of the inverse dependence of the HeMS lifetime ($t_{\mathrm{HeMS}}$) on stellar mass (see Equation 79 in Hurley et al., 2000b). As discussed in Section 4.2, the amount of mass lost due to neutrinos when the VMS transitions into an IMBH is set to 10%, consistent with the default assumptions in NBODY6++GPU [^1] In MOCCA, the default is to assume 1% mass loss due to neutrinos when an evolved star becomes a BH. We also performed an alternative MOCCA run in which the HeMS lifetime of the VMS was extended to 0.3 Myr. In this case, the VMS remained on the HeMS long enough to undergo a collision with a $\rm 53.3\ \leavevmode\nobreak\ \rm M_{\odot}$ stellar-mass BH present in the cluster. Although the mass ratio of this merger was extremely small, if the VMS was allowed to change its stellar type following the collision, the BH was treated as absorbing 50% [^2] The fraction of mass absorbed in collisions between stars and BHs is an input parameter specified in the initial conditions. of the VMS mass. This alternative treatment led to the formation of an IMBH with a mass of $\rm 2.6725\times 10^{4}\ \leavevmode\nobreak\ \rm M_{\odot}$.
@@
Overall, the MOCCA evolution of the VMS mass, radius, and effective age in the simulation presented here is in good agreement with the NBODY6++GPU results, as shown in Fig.  8. The mass distribution and number of stars that collide with the VMS are shown in Fig. 5, where the top right panel highlights that the total number of collisions involving the VMS is lower in MOCCA compared to the direct N-body simulation (top left panel). This results in less efficient rejuvenation, allowing the VMS to reach the end of its main-sequence lifetime and evolve into an IMBH slightly earlier (by about 0.2 Myr) than in the N-body run. Despite the reduced number of collisions, the final VMS mass at the time of IMBH formation is only $\rm 3500\ \leavevmode\nobreak\ \rm M_{\odot}$ smaller in MOCCA, due to the fact that no mass is lost during MS collisions in this simulation.
@@ 5 Summary, conclusions and outlook
This work investigates the formation of VMSs through stellar mergers, using both direct N-body and MOCCA simulations. To better follow the evolution and growth of a VMS in these models, several improvements were made to the standard stellar evolution routines in NBODY6++GPU and MOCCA. These included updates to the treatment of stellar radius evolution, rejuvenation, and close encounter prescriptions. Our simulations demonstrate the growth of a VMS through successive mergers, reaching several tens of thousands of solar masses. The formation of such a massive object through a purely collisional pathway highlights the potential importance of this channel in the formation of IMBHs and SMBH seeds. This dual approach allows us to test the robustness of the VMS formation channel across methods with different computational schemes and resolution limits.
@@ 5.1 Theoretical considerations and simulation constraints
The formation of VMSs through runaway collisions in dense star clusters was first proposed by Begelman & Rees (1978), Rees (1984), and Lee (1987) using simple analytical models. This scenario was later explored using Fokker–Planck and Monte Carlo methods by Quinlan & Shapiro (1990) and Freitag et al. (2006d, b), and more recently through direct N-body simulations by Vergara et al. (2023). In clusters with a larger number of particles, dynamical equilibrium is typically reached before thermodynamic equilibrium. As a result, such systems require longer times to undergo core collapse and reach Spitzer instability without significant expansion due to relaxation processes (Spitzer, 1987). Our results support the existence of a critical mass threshold for the formation of massive objects via collisions, as previously suggested by Vergara et al. (2023, 2024). Our simulations push into an uncharted regime of extremely high central density and particle number, testing the runaway collision scenario under extreme but plausible cluster conditions.
@@
The high central density of the clusters in our models ($>10^{7}\leavevmode\nobreak\ \rm M_{\odot}\leavevmode\nobreak\ \rm pc^{-3}$) leads to an elevated stellar collision rate, significantly higher than that in typical observed globular clusters. Consequently, the time between successive collisions becomes much shorter than the thermal timescale of the VMS. This rapid succession of collisions prevents the VMS from reaching thermal equilibrium between impacts, potentially causing it to expand beyond its equilibrium radius, which may, in turn, alter its collision cross-section and affect the likelihood and nature of subsequent encounters. The physical and thermodynamical structure of such a VMS lies in a largely unexplored regime, and it may experience significant mass loss through repeated collisions and strong stellar winds, both of which can affect its growth in radius and mass (Dale & Davies, 2006; Suzuki et al., 2007). As a result, our simulation likely provides an optimistic upper limit on the achievable VMS mass under such conditions.
@@
Repeated collisions also rejuvenate the VMS, delaying its natural evolution toward collapse into a BH. VMSs may be important contributors to cosmic hydrogen and helium reionization, a process that depends sensitively on stellar wind mass-loss rates and metallicity (Vink et al., 2021; Sander et al., 2020). In our models, we adopt a relatively high metallicity of $Z=0.01$, which enhances wind-driven mass loss and makes VMS formation more difficult compared to lower-metallicity environments. Despite this, the VMS forms rapidly enough that winds do not inhibit its growth, owing to the assumption that the star returns to thermal equilibrium following each collision. Between the two codes, the higher collision rate in the N-body run leads to more frequent rejuvenation and prolongs the main-sequence lifetime of the VMS compared to the MOCCA run. The stellar wind prescriptions in both simulations are based on the works of Vink et al. (2001), Vink & de Koter (2002, 2005), and Belczynski et al. (2010). As the VMS evolves into the helium-burning phase, it experiences stronger but shorter-lived wind episodes. For this phase, we adopt a luminosity-dependent wind recipe following Sander et al. (2020). However, since the helium main-sequence lifetime is relatively short ($\sim 0.01\leavevmode\nobreak\ \rm Myr$) for the VMS in these runs, there is no significant mass loss due to winds during this evolutionary stage. In future work, we plan to implement a more refined mass–luminosity relation to improve the modeling of VMS evolution and feedback.
@@
The early and rapid mass gain in our models is essential; had the VMS grown more slowly, stellar winds at $Z=0.01$ would likely have halted its development, preventing IMBH formation. However, this outcome also depends on how strongly the VMS deviates from thermal equilibrium after collisions, as larger radii could enhance the stellar collision rate and potentially compensate for wind-driven mass loss. Furthermore, the IMBH forms through the direct collapse of the VMS, with an assumed mass loss of approximately 10% due to neutrino emission (Fryer et al., 2012; Kamlah et al., 2022a). However, the exact amount of mass lost, if any, during the transition from an evolved star to a BH remains uncertain.
@@
VMS formation at high metallicities ($Z=0.01$–0.02) has also been investigated using smoothed-particle hydrodynamics (SPH) simulations with ASURA-BRIDGE (Fujii et al., 2024a). These studies model dense clusters with central densities of $10^{6}$$10^{9}\leavevmode\nobreak\ \rm M_{\odot}\,pc^{-3}$ and show that IMBHs with masses exceeding $4000\leavevmode\nobreak\ \rm M_{\odot}$ can form within approximately 1.5 Myr (Fujii et al., 2024b). The inclusion of gas in these models contributes to the higher central densities and enhances the conditions for runaway collisions and rapid VMS growth. Unlike our gas-free N-body and Monte Carlo models, these simulations demonstrate how gas inflows can further support the VMS formation channel. Nevertheless, our results show that even in purely stellar systems, rapid VMS formation is feasible under extreme initial conditions.
@@
If we interpret runaway stellar collisions as analogous to gas accretion, we can compare our results to models of accreting stars (Haemmerlé et al., 2018; Herrington et al., 2023). In particular, we focus on models with an accretion rate of $\sim 0.01\leavevmode\nobreak\ \rm M_{\odot}\,yr^{-1}$, which is comparable to the effective rate at which our VMS gains mass through mergers. These models predict that accreting stars expand to radii of $\sim 10^{4}\leavevmode\nobreak\ \rm R_{\odot}$ along the Hayashi track. Notably, Herrington et al. (2023) terminate their simulations at central hydrogen depletion. The VMS in our simulations reaches a radius comparable to those found in these accreting-star models. Since stellar radius is an important parameter influencing collision rates and outcomes, the agreement between our VMS radii and those predicted in accretion models supports the plausibility of the VMS growth we observe. However, the internal structure of a VMS built through collisions may differ significantly from that of an accreting MS star. In particular, collisional VMSs may develop a dense core with an extended, dilute envelope, potentially reducing the efficiency of further mass buildup by lowering the effective collision cross-section. This structural difference adds uncertainty when interpreting the final VMS mass in our simulations.
@@
Furthermore, Martinet et al. (2023) present stellar evolution models computed with the GENEC code (Eggenberger et al., 2008) for metallicity $Z=0.014$, including both rotating and non-rotating stars with initial masses of 180, 250, and $300\leavevmode\nobreak\ \rm M_{\odot}$. These models yield main-sequence lifetimes of approximately 2.2–2.6 Myr. In comparison, the SSE models by Hurley et al. (2000a), which we use in our simulations, give a main-sequence lifetime of 2.9 Myr at $Z=0.01$. This agreement suggests that the stellar evolution prescriptions adopted in our simulations are consistent with more detailed stellar evolution models, lending credibility to our estimates of VMS lifetimes and their impact on subsequent cluster evolution.
@@
The radiative efficiency of stellar winds from VMSs could play an important role in the chemical enrichment of star clusters. This process has been proposed as a potential mechanism for the formation of multiple stellar populations (MSPs) (Gieles et al., 2018, 2025). Recent high-resolution hydrodynamic simulations have provided the first detailed insights into VMS self-enrichment in massive clusters, highlighting its relevance for MSP formation. Stellar winds enriched with elements such as Na, Al, and N can be retained within the cluster and contribute to the formation of chemically distinct stars within just a few million years (Lahén et al., 2024). In addition to stellar winds, stellar mergers offer an alternative pathway for MSP formation. In dense clusters where mergers are more frequent, these events can eject chemically enriched material into the surrounding gas. This material—whether expelled through dynamical interactions, explosions, or fusion-driven outflows—can lead to the formation of new stars with distinct chemical compositions (Wang et al., 2020b). Recent observational studies have also revealed detailed kinematic differences between stellar populations, showing that second-generation stars tend to rotate more rapidly and are more centrally concentrated than first-generation stars (Dalessandro et al., 2024). These findings provide further motivation for detailed modeling of early cluster enrichment scenarios involving VMSs, especially in young massive clusters forming at high redshift.

### v2

@@ 4.2 N-body results: VMS evolution and IMBH formation
The hydrogen burning phase is set to end in 2.92.9\,Myr, according to the SSE/BSE prescription Hurley2000; Hurley2002a; Hurley2005. After the VMS reaches this point, it should start burning helium. At 4.674.67\,Myr the VMS has a mass of $5.7362\times 10^{4}\penalty 10000\ \rm M_{\odot}$, with a radius of $1.9\times 10^{5}\rm R_{\odot}$ and effective age of 2.912.91\,Myr. At this stage, the VMS becomes a helium main sequence (HeMS; or type 7 in SSE/BSE by Hurley2000; Hurley2002a; Hurley2005)) star with a radius of $171\rm R_{\odot}$. During this helium phase, the VMS experiences three collisions with MS stars within 0.020.02\,Myr before naturally collapsing into an IMBH. This process may result in mass loss of about $10\penalty 10000\ \%$ due to neutrino emission (Fryer2012; Kamlah2022a), leading to a mass of $51\,625.99\penalty 10000\ \rm M_{\odot}$ at time 4.694.69\,Myr. Finally, the IMBH undergoes additional collisions, accreting $50\%$ of the stellar companion (Rizzuto2021), reaching a final mass of $51\,661.21\penalty 10000\ \rm M_{\odot}$ by the end of the simulation.
@@ 4.3 MOCCA results: VMS evolution and IMBH formation
In the MOCCA simulation, the star that eventually grows into a VMS has an initial mass of $\rm 130.6\ \penalty 10000\ \rm M_{\odot}$ and is initially located at a distance of 0.088 pc from the center of the cluster. Within 0.15 Myr, this massive star undergoes four collisions with low mass stars, slightly increasing its mass to approximately $\rm 132\ \penalty 10000\ \rm M_{\odot}$. For a brief period between about $0.18$ Myr and $0.21$ Myr, the VMS is part of a binary system. At $0.182$ Myr, the VMS first forms a binary with a $\rm 72.4\,M_{\odot}$ star due to 3BBF. Shortly after, at $0.183$ Myr, this companion is replaced by a $\rm 124\,M_{\odot}$ star in an exchange encounter. Following several flyby interactions, the VMS and its binary companion merge during a binary-single interaction at $0.186$ Myr. The VMS subsequently forms another binary, first with a $\rm 72.4\,M_{\odot}$ star, then with a $\rm 68\,M_{\odot}$ star, and later with a $\rm 174\,M_{\odot}$ star. The latter also merges with the VMS in a binary-single interaction at around $0.21$ Myr, following several flyby encounters. At this stage, the VMS is located at a distance of approximately $3\times 10^{-4}\ \rm pc$ from the cluster center. By $0.212$ Myr, its mass exceeds $\rm 500\,M_{\odot}$ following a collision with a $\rm 121.8\,M_{\odot}$ star, with which it was also briefly in a binary system. As described in Section 3.2, from this point onward, we disable 3BBF for the VMS. Its mass continues to increase rapidly as it undergoes runaway collisions with other MS stars in the cluster center. By 0.23 Myr, the VMS mass reaches $\rm 1000\ \penalty 10000\ \rm M_{\odot}$, and by 0.34 Myr, it exceeds $\rm 5000\ \penalty 10000\ \rm M_{\odot}$. The growth of the VMS in MOCCA is shown in the top-right panel of Fig. 8.
@@
During these early stages of VMS growth, collisions with other VMSs also strongly rejuvenate it. At 0.34 Myr, when its mass reaches $\rm 5000\ \penalty 10000\ \rm M_{\odot}$, its effective age is only 0.068 Myr, and its radius is approximately $\rm 68\ \mathrm{R}_{\odot}$. By 0.51 Myr, the VMS mass surpasses $\rm 10^{4}\ \penalty 10000\ \rm M_{\odot}$, at which point its effective age is 0.17 Myr, and its radius is $\rm 71.5\ \mathrm{R}_{\odot}$. The VMS continues its rapid growth, reaching $\rm 2.1\times 10^{4}\ \penalty 10000\ \rm M_{\odot}$ by 1 Myr. By this time, the cumulative number of collisions involving the VMS is approximately 3,815 (see Fig. 3). As the mass ratio of these collisions decreases with increasing VMS mass, its effective rejuvenation per collision also weakens. At 1 Myr, its effective age is 0.468 Myr. The evolution of the VMS’s effective age on the MS is shown in the bottom-right panel of Fig. 8.
@@
By 2 Myr, the VMS reaches a mass of $\rm 3.56\times 10^{4}\ \penalty 10000\ \rm M_{\odot}$, with an effective age of 1.1 Myr and a radius of $\rm 131\ \mathrm{R}_{\odot}$. Up until this point, the VMS has undergone more than 6,793 collisions. Between 2 Myr and 3 Myr, its growth slows, and by 3 Myr, it reaches a mass of $\rm 4.59\times 10^{4}\ \penalty 10000\ \rm M_{\odot}$ with an effective age of 1.77 Myr. This corresponds to approximately 61% of the MS lifetime of the VMS, which is set to 2.91 Myr in the SSE/BSE prescriptions used in NBODY6++GPU and MOCCA. As the VMS ages, its radius increases, reaching $\rm 523\ \mathrm{R}_{\odot}$ at 3 Myr. The evolution of the VMS’s radius is shown in the middle-right panel of Fig. 8.
@@
At 4.477 Myr, the VMS reaches a mass of $\rm 5.38\times 10^{4}\ \penalty 10000\ \rm M_{\odot}$, with a radius of $\rm 1.9\times 10^{5}\mathrm{R}_{\odot}$ and an effective age of 2.91 Myr, marking the MS turn-off time for the VMS. By this point, it has undergone 10,322 collisions. After this stage, the VMS evolves into a helium main sequence star (HeMS; type 7 in SSE/BSE) with a radius of $\rm 165\ \mathrm{R}_{\odot}$. Between 4.477 Myr and 4.485 Myr, the HeMS VMS undergoes approximately 15 additional collisions with lower-mass stars, reaching a final mass of $\rm 5.384\times 10^{4}\ \penalty 10000\ \rm M_{\odot}$. As the mass ratios in these collisions are extremely small, no rejuvenation is applied. At 4.485 Myr, the VMS evolves into an IMBH with a mass of $\rm 4.845\times 10^{4}\ \penalty 10000\ \rm M_{\odot}$. This IMBH then continues to grow through collisions, mostly with other stars, assuming that 50% of the stellar mass is accreted by the IMBH in each collision. By 5 Myr, it reaches a mass of $\rm 4.984\times 10^{4}\ \penalty 10000\ \rm M_{\odot}$.
@@
The short duration of the HeMS phase in the evolution of the VMS is a result of the inverse dependence of the HeMS lifetime ($t_{\mathrm{HeMS}}$) on stellar mass (see Equation 79 in Hurley2000). As discussed in Section 4.2, the amount of mass lost due to neutrinos when the VMS transitions into an IMBH is set to 10%, consistent with the default assumptions in NBODY6++GPU [^1] In MOCCA, the default is to assume 1% mass loss due to neutrinos when an evolved star becomes a BH. We also performed an alternative MOCCA run in which the HeMS lifetime of the VMS was extended to 0.3 Myr. In this case, the VMS remained on the HeMS long enough to undergo a collision with a $\rm 53.3\ \penalty 10000\ \rm M_{\odot}$ stellar-mass BH present in the cluster. Although the mass ratio of this merger was extremely small, if the VMS was allowed to change its stellar type following the collision, the BH was treated as absorbing 50% [^2] The fraction of mass absorbed in collisions between stars and BHs is an input parameter specified in the initial conditions. of the VMS mass. This alternative treatment led to the formation of an IMBH with a mass of $\rm 2.6725\times 10^{4}\ \penalty 10000\ \rm M_{\odot}$.
@@
Overall, the MOCCA evolution of the VMS mass, radius, and effective age in the simulation presented here is in good agreement with the NBODY6++GPU results, as shown in Fig.  8. The mass distribution and number of stars that collide with the VMS are shown in Fig. 5, where the top right panel highlights that the total number of collisions involving the VMS is lower in MOCCA compared to the direct N-body simulation (top left panel). This results in less efficient rejuvenation, allowing the VMS to reach the end of its main-sequence lifetime and evolve into an IMBH slightly earlier (by about 0.2 Myr) than in the N-body run. Despite the reduced number of collisions, the final VMS mass at the time of IMBH formation is only $\rm 3500\ \penalty 10000\ \rm M_{\odot}$ smaller in MOCCA, due to the fact that no mass is lost during MS collisions in this simulation.
@@ 5 Summary, conclusions and outlook
This work investigates the formation of VMSs through stellar mergers, using both direct N-body and MOCCA simulations. To better follow the evolution and growth of a VMS in these models, several improvements were made to the standard stellar evolution routines in NBODY6++GPU and MOCCA. These included updates to the treatment of stellar radius evolution, rejuvenation, and close encounter prescriptions. We note that the prescriptions implemented here represent reasonable extrapolations of standard SSE/BSE stellar evolution models (Hurley2000; Hurley2002a; Hurley2005; Kamlah2022a); however, the internal structure and thermal state of VMS formed through stellar collisions remain to be established and require further refinement. Our simulations demonstrate the growth of a VMS through successive mergers, reaching several tens of thousands of solar masses. The formation of such a massive object through a purely collisional pathway highlights the potential importance of this channel in the formation of IMBHs and SMBH seeds. This dual approach allows us to test the robustness of the VMS formation channel across methods with different computational schemes and resolution limits. We note that the MOCCA results required some adjustments to achieve consistency with the NBODY6++GPU simulations. In particular, the timestep was reduced to better capture the mass segregation time of massive stars. Without this change, a VMS would still have formed and reached a comparable final mass to the N-body run, but its growth would be delayed by $\approx$0.2 Myr. In addition, 3BBF was disabled for the VMS once it exceeded $500\penalty 10000\ \rm M_{\odot}$. With 3BBF enabled, the VMS quickly forms a binary and is prevented from undergoing two-body collisions, reaching only $\sim 35{,}000\penalty 10000\ \rm M_{\odot}$ at $4.5$ Myr. In contrast, with 3BBF disabled, the collisional growth of the VMS proceeds at a rate consistent with the direct N-body result. These points are discussed in more detail in Subsection 3.2.
@@ 5.1 Theoretical considerations and simulation constraints
The formation of VMSs through runaway collisions in dense star clusters was first proposed by Begelman1978, Rees1984, and Lee1987 using simple analytical models. This scenario was later explored using Fokker–Planck and Monte Carlo methods by Quinlan1990; Gurkan2004; Freitag2006a; Freitag2006 and Sharma2025, and more recently through direct N-body simulations by Vergara2023. In clusters with a larger number of particles, dynamical equilibrium is typically reached before thermodynamic equilibrium. As a result, such systems require longer times to undergo core collapse and reach Spitzer instability without significant expansion due to relaxation processes (Spitzer1987). Our results support the existence of a critical mass threshold for the formation of massive objects via collisions, as previously suggested by Vergara2023; Vergara2024. Our simulations push into an uncharted regime of extremely high central density and particle number, testing the runaway collision scenario under extreme but plausible cluster conditions.
@@
The high central density of the clusters in our models ($10^{7}\penalty 10000\ \rm M_{\odot}\penalty 10000\ \rm pc^{-3}$) leads to an elevated stellar collision rate, significantly higher than that in typical observed globular clusters. Consequently, the time between successive collisions becomes much shorter than the thermal timescale of the VMS. This rapid succession of collisions prevents the VMS from reaching thermal equilibrium between impacts, potentially causing it to expand beyond its equilibrium radius, which may, in turn, alter its collision cross-section and affect the likelihood and nature of subsequent encounters. The physical and thermodynamical structure of such a VMS lies in a largely unexplored regime, and it may experience significant mass loss through repeated collisions and strong stellar winds, both of which can affect its growth in radius and mass (Dale2006; Suzuki2007). As a result, our simulation likely provides an optimistic upper limit on the achievable VMS mass under such conditions.
@@
Repeated collisions also rejuvenate the VMS, delaying its natural evolution toward collapse into a BH. VMSs may be important contributors to cosmic hydrogen and helium reionization, a process that depends sensitively on stellar wind mass-loss rates and metallicity (Vink2021; Sander2020). In our models, we adopt a relatively high metallicity of $Z=0.01$, which enhances wind-driven mass loss and makes VMS formation more difficult compared to lower-metallicity environments. Despite this, the VMS forms rapidly enough that winds do not inhibit its growth, owing to the assumption that the star returns to thermal equilibrium following each collision. Between the two codes, the higher collision rate in the N-body run leads to more frequent rejuvenation and prolongs the main-sequence lifetime of the VMS compared to the MOCCA run. The stellar wind prescriptions in both simulations are based on the works of Vink2001, Vink2002; Vink2005, and Belczynski2010. As the VMS evolves into the helium-burning phase, it experiences stronger but shorter-lived wind episodes. For this phase, we adopt a luminosity-dependent wind recipe following Sander2020. However, since the helium main-sequence lifetime is relatively short ($\sim 0.01\penalty 10000\ \rm Myr$) for the VMS in these runs, there is no significant mass loss due to winds during this evolutionary stage. In future work, we plan to implement a more refined mass–luminosity relation to improve the modeling of VMS evolution and feedback.
@@
The early and rapid mass gain in our models is essential; had the VMS grown more slowly, stellar winds at $Z=0.01$ would likely have halted its development, preventing IMBH formation. However, this outcome also depends on how strongly the VMS deviates from thermal equilibrium after collisions, as larger radii could enhance the stellar collision rate and potentially compensate for wind-driven mass loss. Furthermore, the IMBH forms through the direct collapse of the VMS, with an assumed mass loss of approximately 10% due to neutrino emission (Fryer2012; Kamlah2022a). However, the exact amount of mass lost, if any, during the transition from an evolved star to a BH remains uncertain.
@@
VMS formation at high mlicities (Z=0.01Z=0.01–0.02) has also been investigated using smoothed-particle hydrodynamics (SPH) simulations with ASURA-BRIDGE (Fujii2024a). These studies model dense clusters with central densities of $10^{6}$$10^{9}\penalty 10000\ \rm M_{\odot}\,pc^{-3}$ and show that IMBHs with masses exceeding $4000\penalty 10000\ \rm M_{\odot}$ can form within approximately 1.5 Myr (Fujii2024). The inclusion of gas in these models contributes to the higher central densities and enhances the conditions for runaway collisions and rapid VMS growth. Unlike our gas-free N-body and Monte Carlo models, these simulations demonstrate how gas inflows can further support the VMS formation channel. Nevertheless, our results show that even in purely stellar systems, rapid VMS formation is feasible under extreme initial conditions.
@@
If we interpret runaway stellar collisions as analogous to gas accretion, we can compare our results to models of accreting stars (Haemmerle2018; herrington2023). In particular, we focus on models with an accretion rate of $\sim 0.01\penalty 10000\ \rm M_{\odot}\,yr^{-1}$, which is comparable to the effective rate at which our VMS gains mass through mergers. These models predict that accreting stars expand to radii of $\sim 10^{4}\penalty 10000\ \rm R_{\odot}$ along the Hayashi track. Notably, herrington2023 terminate their simulations at central hydrogen depletion. The VMS in our simulations reaches a radius comparable to those found in these accreting-star models. Since stellar radius is an important parameter influencing collision rates and outcomes, the agreement between our VMS radii and those predicted in accretion models supports the plausibility of the VMS growth we observe. However, the internal structure of a VMS built through collisions may differ significantly from that of an accreting MS star. In particular, collisional VMSs may develop a dense core with an extended, dilute envelope, potentially reducing the efficiency of further mass buildup by lowering the effective collision cross-section. This structural difference adds uncertainty when interpreting the final VMS mass in our simulations.
@@
Furthermore, Martinet2023 present stellar evolution models computed with the GENEC code (Eggenberger2008) for metallicity $Z=0.014$, including both rotating and non-rotating stars with initial masses of 180, 250, and $300\penalty 10000\ \rm M_{\odot}$. These models yield main-sequence lifetimes of approximately 2.2–2.6 Myr. In comparison, the SSE models by Hurley2000, which we use in our simulations, give a main-sequence lifetime of 2.9 Myr at $Z=0.01$. This agreement suggests that the stellar evolution prescriptions adopted in our simulations are consistent with more detailed stellar evolution models, lending credibility to our estimates of VMS lifetimes and their impact on subsequent cluster evolution.
@@
The radiative efficiency of stellar winds from VMSs could play an important role in the chemical enrichment of star clusters. This process has been proposed as a potential mechanism for the formation of multiple stellar populations (MSPs) (Gieles2018; Gieles2025). Recent high-resolution hydrodynamic simulations have provided the first detailed insights into VMS self-enrichment in massive clusters, highlighting its relevance for MSP formation. Stellar winds enriched with elements such as Na, Al, and N can be retained within the cluster and contribute to the formation of chemically distinct stars within just a few million years (Lahe2024). In addition to stellar winds, stellar mergers offer an alternative pathway for MSP formation. In dense clusters where mergers are more frequent, these events can eject chemically enriched material into the surrounding gas. This material—whether expelled through dynamical interactions, explosions, or fusion-driven outflows—can lead to the formation of new stars with distinct chemical compositions (Wang2020e). Recent observational studies have also revealed detailed kinematic differences between stellar populations, showing that second-generation stars tend to rotate more rapidly and are more centrally concentrated than first-generation stars (Dalessandro2024). These findings provide further motivation for detailed modeling of early cluster enrichment scenarios involving VMSs, especially in young massive clusters forming at high redshift.